OpenStudy (gokuporter):

Which sets of numbers are closed under addition? Choose all answers that are correct. A. whole numbers B. natural numbers C. negative integers D. integers

OpenStudy (perl):

closure under an operation means that if you start with two elements in a set and then operate on them, the result will also be in the set

OpenStudy (perl):

any two elements

OpenStudy (perl):

one sec, draw button died on me

OpenStudy (perl):



OpenStudy (perl):

well 1 + 1 = 2 , and 2 is inside the set . 1+ 2 = 3 and 3 is inside the set .... looks like it is closed

OpenStudy (perl):

0 + 1 = 1 , 0 + 0 = 0 , etc

OpenStudy (perl):

how about a set that is not closed under an operation. lets look at natural numbers under subtraction . is that closed?

OpenStudy (perl):

, well if we pick two numbers from the set of natural numbers, how about 1 and 2. 1 - 2 = -1 but -1 is not a natural number. so the natural numbers are not closed under subtraction

OpenStudy (perl):

so in your problem, A through D are closed under addition.

OpenStudy (perl):

anybody home?

OpenStudy (gokuporter):

Oh hey whats up